The difference of a number m and twice a number n is 1. The sum of two times m and n is 22. I believe the system of equation would be:
m-2n=1
2(m+n)=22
m-2n=1
2m+2n=22 or is it
m-2n=1
2m+n=22. Ugh, I am really need guidance how to write the second equation. Many thanks.
AC3

Answer by Cromlix(4381)

m - 2n = 1.....(1)
2m + n = 22.....(2)
Multiply (2) by 2
m - 2n = 1......(1)
4m + 2n = 44.....(3)
Add (1) + (3)
5m = 45
m = 9
Substitute m = 9 into (1)
m - 2n = 1......(1)
9 - 2n = 1
-2n = 1 - 9
-2n = -8
n = 4
m = 9 and n = 4.
Hope this helps.
:-)